10. Which of the functions corresponds to the graph? 10+ 10. A f(x)=6+3 B. f(x)=e* +4 C. f(x)=et +4 D. f(x)=e*+3
Answer #1

From the graph

At x = 0 f(x) =4

i.e f(0)=4

On checking the given options

A) f(0)= -e^0+3 = -1+3 =2

B)f(0)= e^0+4 = 1+4=5

C)f(0)=e^0+4= 1+4=5

D)f(0)=e^0+3 = 1+3 = 4

So, the answer is option D

i.e f(x)=e^-x + 3